OpenStudy (anonymous):

an online coupon service has 8 offers for free samples. how many different requests are possible if A: a customer must request EXACTLY 2 free samples. B: a customer must request UP TO 2 free samples.

OpenStudy (kropot72):

A: The number of combinations of 8 offers taken 2 at a time is\[8C2=\frac{8!}{6!2!}=\frac{8\times 7}{2}\] B: Multiply the answer to A: by 8.

OpenStudy (anonymous):

B is wrong :/

OpenStudy (kropot72):

B: Try adding 8 to the answer to B:

OpenStudy (anonymous):

nooooo ;(

OpenStudy (kropot72):

Maybe I don't understand the question. There are 8 ways of using the offers taken 1 at a time. There are 28 ways of using the offers taken two at a time.